OpenStudy (anonymous):

sin(2x) divided by 6 equals?

OpenStudy (anonymous):

use a calculator thats provided online

terenzreignz (terenzreignz):

Use this identity \[\LARGE \sin(2x) = 2\sin(x)\cos(x)\]

OpenStudy (anonymous):

thanks